# Nature Language Model: Deciphering the Language of Nature for Scientific Discovery

### v2

@@ 4 Protein tasks
    Protein description generation: This task focuses on generating explanations or uncovering properties and functions of protein sequences, articulating them in natural language.
@@ 4.1 Unconditioned generation
Mixtral 8x7b 53.3 0.906 69.9
GPT-4 45.7 0.816 65.1
NatureLM (1B) 288.3 0.985 69.8
NatureLM (8B) 284.5 0.973 71.8
NatureLM (8x7B) 318.4 0.989 75.9
Table 8: Protein Sequence Generation Comparison. The average length of natural proteins (calculated from a subset of proteins randomly sampled from UR50) is about 311. The diversity was calculated by the number of clusters with 50% sequence identity divided by the total generated sequence count. The pLDDT scores were calculated by OmegaFold omegafold on the generated sequences with length less than 100 for a fair comparison. The length distribution is left in Figure S16.
As shown in Table 8, NatureLM consistently outperform Mixtral 8x7b and GPT-4 in terms of average sequence length, diversity, and average pLDDT score. The NatureLM (8x7B) model achieves the best performance across all metrics, with an average length of 318.4, diversity of 0.989, and average pLDDT score of 75.9. ProLLAMA lv2024prollamaproteinlanguagemodel a fine-tuned LLM for protein. It generates proteins without explicitly defined constraints on length, achieving a pLDDT score of 66.5. In contrast, our approach, which does not impose length constraints, results in pLDDT scores of 69.8 and 78.1 for the 8B and 8x7B models, respectively, demonstrating our significant advancement in this area.
@@ 4.2 Text-guided protein generation
For text-guided protein generation, we evaluated our models’ ability to generate proteins with specific properties based on natural language prompts. In this study, we focused on two key properties: solubility and stability, leaving the exploration of additional properties for future work. For stability, the models were tasked with generating protein sequences that exhibit stable properties. Regarding solubility, since both soluble and insoluble proteins are common in natural sequences, we instructed NatureLM to generate sequences of both types. Sample prompts are shown below, and a full list of prompts can be found in Figure S18.
@@ 4.3 Antigen-binding CDR-H3 design
Table 10 presents the Amino Acid Recovery (AAR) scores for the CDR-H3 design task. As the model size of NatureLM increase, the AAR gradually increases. The NatureLM (8x7B) model achieves competitive performance with an AAR of 0.376, outperforming several specialized GNN-based models. While SFM-protein, a BERT-like model trained on protein sequences, holds the top performance, our results demonstrate the potential of NatureLM in CDR-H3 design, particularly as the model scales and undergoes further refinement.
4.4 Protein description generation
@@
To achieve this, we compiled pairs of protein sequences and their human-readable annotations from various species, sourced from the NCBI database. We divided the dataset temporally: historical data were utilized for training the NatureLM models, while annotation data from the most recent four months were reserved for testing. Model performance was evaluated using Rouge-L scores. As shown in Table 11, NatureLM models consistently outperformed Llama 3 8B in Rouge-L scores, with performance differences widening as model size increased. Notably, the NatureLM (8x7B) model achieved the highest score of 0.585. A detailed analysis presented in Figure 16 revealed that the NatureLM (8x7B) model not only generates protein annotations with greater accuracy but also successfully identifies orthologues and functions of proteins, while NatureLM (8B) is also able to generate reasonable results in many cases.
@@ 4.5 Heme-binding protein design driven by text and SMILES
Step 4: structure generation and validation. For each protein in HemeProtein, we use Protenix Protenix2025 to predict the complex structure of the generated proteins bound to heme. The predicted structures are then inspected to ensure that the proteins can form the critical interaction with heme for stable binding.
@@
In the text-based design case study (Fig.17 left), two histidine residues are positioned in close proximity to the iron located in the center of heme, enabling the formation of coordinated bonding interactions with the heme group. Similarly, in the SMILES-based design, the algorithm can output proteins with binding motifs similar to those generated in the text-based example (Fig. S13). However, as shown in Fig. 17 (right), we show a representative case where a methionine and histidine residue are observed to interact closely with the iron ion (see Fig. S14 for more discussion on this case). These residues effectively coordinate the metal ion through their respective side chains, demonstrating alternative structural strategies for heme binding. Furthermore, the designed protein sequences differ significantly from those present in the database, indicating that our approach can generate novel sequences with distinct structural properties. Collectively, these results demonstrate the effectiveness of NatureLM in designing functional heme-binding proteins with diverse and novel structural features. We also compare the apo and holo structures of the generated proteins in Fig. S15, which shows that the key residues involved in heme binding, such as histidine and methionine, occupy similar positions in both structures.
@@ 5.1 Unconditional material generation
5.2 Composition to material generation
@@
5.3 Bulk modulus to material generation
The bulk modulus of a substance is a measure of the resistance of a substance to bulk compression. As a proof-of-concept, the model is prompted to generate materials with specified bulk modulus:
@@ 5.3 Bulk modulus to material generation
We evaluated the SMACT validity, stability, novelty, and precision of the generated materials. Precision is defined as the ratio of generated materials whose bulk modulus is within 10% of the instructed value, compared to all generated materials.
@@ 6.2 Guide RNA design
Guided RNA, commonly referred to as guide RNA (gRNA), is a key element in CRISPR-Cas9 gene-editing technology. It is essential for directing the Cas9 enzyme to a precise location within the genome where genetic modifications are intended. We evaluate NatureLM on two gRNA design tasks: the first is designing gRNAs for a given DNA sequence, and the second is selecting the more effective gRNA from two candidates. Examples are provided below:
@@
A valid gRNA is defined by the following criteria: (1) the length of the generated gRNA must be between 17 to 24 nucleotides; (2) the generated sequences should match a region in the given DNA sequences; (3) the matched sequence in the DNA template must be directly followed by an “NGG” sequence, known as the Protospacer Adjacent Motif (PAM).
@@ Appendix A Supplementary figures
Figure S1: Correct ratio of property-to-molecule generation. We treat the generated molecule as a correct one if $\lvert v^{\prime}-v\lvert\leq\delta$, where $v^{\prime}$ is its property value and $v$ is the input value. $\delta$ is set to 0 for HBA, HBD, RotBonds, 0.05 for QED and FSP3, and 5 for TPSA.

### v3

@@ 4 Protein tasks
    Protein characteristics description generation: This task is to generate clear, human-readable explanations of protein sequences, describing their properties and functions.
  5. 5.
    Case study: The model designs heme-binding protein design driven by text and SMILES.
@@ 4.1 Unconditioned generation
Mixtral 8x7b 53 0.906 69.9
GPT-4 46 0.816 65.1
NatureLM (1B) 288 0.985 69.8
NatureLM (8B) 285 0.973 71.8
NatureLM (8x7B) 318 0.989 75.9
Table 8: Protein Sequence Generation Comparison. The average length of natural proteins (calculated from a subset of proteins randomly sampled from UR50) is about 311. The diversity was calculated by the number of clusters with 50% sequence identity divided by the total generated sequence count. The pLDDT scores were calculated by OmegaFold omegafold on the generated sequences with length less than 100 for a fair comparison. The length distribution is shown in Figure S16.
As shown in Table 8, NatureLM consistently outperform Mixtral 8x7b and GPT-4 in terms of average sequence length, diversity, and average pLDDT score. The NatureLM (8x7B) model achieves the best performance across all metrics, with an average length of 318, diversity of 0.989, and average pLDDT score of 75.9. ProLLAMA lv2024prollamaproteinlanguagemodel , a fine-tuned LLM for protein, generates proteins without explicitly defined constraints on length, achieving a pLDDT score of 66.5. In contrast, our approach, which does not impose length constraints, results in pLDDT scores of 69.8 and 78.1 for the 8B and 8x7B models, respectively, demonstrating our significant advancement in protein sequence generation.
@@ 4.2 Text-guided protein generation
For text-guided protein generation, we evaluated our models’ ability to generate proteins with specific properties based on natural language prompts. In this study, we focused on two key properties: solubility and stability, leaving the exploration of additional properties for future work. For stability, the models were tasked with generating protein sequences that exhibit stable properties. Regarding solubility, since both soluble and insoluble proteins exist in nature, we instructed NatureLM to generate sequences of both types. Sample prompts are shown below, and a full list of prompts can be found in Figure S18.
@@ 4.3 Antigen-binding CDR-H3 design
Table 10 presents the Amino Acid Recovery (AAR) scores for the CDR-H3 design task. As the model size of NatureLM increases, the AAR gradually increases. The NatureLM (8x7B) model achieves competitive performance with an AAR of 0.376, outperforming several specialized GNN-based models, except ABGNN gao2023abgnn in our study. While SFM-protein, a BERT-like model trained on protein sequences, holds the top performance, our results demonstrate the potential of NatureLM in CDR-H3 design, particularly as the model scales and undergoes further refinement.
4.4 Protein characteristics description generation
@@
To achieve this, we compiled pairs of protein sequences and their human-readable annotations from various species, sourced from the NCBI database. We divided the dataset temporally. Model performance was evaluated using Rouge-L scores. As shown in Table 11, NatureLM models consistently outperformed Llama 3 8B in Rouge-L scores, with performance differences widening as model size increased. Notably, the NatureLM (8x7B) model achieved the highest score of 0.585. A detailed analysis presented in Figure 16 revealed that the NatureLM (8x7B) model not only generates protein annotations with higher accuracy but also successfully identifies orthologues and functions of proteins, while NatureLM (8B) is also able to generate reasonable results in many cases.
@@ 4.5 Heme-binding protein design driven by text and SMILES
Step 4: structure prediction and validation. For each protein in HemeProtein, we use Protenix Protenix2025 to predict the complex structure of the generated proteins bound to heme. The predicted structures are then inspected to ensure that the proteins can form the critical interaction with heme for stable binding.
@@
In the text-based design case study (Fig.17 left), two histidine residues are positioned in close proximity to the iron located in the center of heme, enabling the formation of coordinated bonding interactions with the heme group. Similarly, in the SMILES-based design, the NatureLM generates proteins with binding motifs similar to those generated in the text-based example (Fig. S13). However, as shown in Fig. 17 (right), we show a representative case where a methionine and histidine residue are observed to interact closely with the iron ion (see Fig. S14 for more discussion on this case). These residues effectively coordinate the metal ion through their respective side chains, demonstrating alternative structural strategies for heme binding. Furthermore, the designed protein sequences differ significantly from those present in the database, indicating that NatureLM can generate novel sequences with distinct structural properties. Collectively, these results demonstrate the effectiveness of NatureLM in designing functional heme-binding proteins with diverse and novel structural features. We also compare the apo and holo structures of the generated proteins in Fig. S15, which shows that the key residues involved in heme binding, such as histidine and methionine, occupy similar positions in both structures.
@@ 5.1 Unconditional material generation
5.2 Material generation for given composition
@@
5.3 Material generation for desired properties
NatureLM can generate materials with desired properties, and here we demonstrate this capability by generating materials that have specific bulk modulus. The bulk modulus of a substance is a measure of the resistance of a substance to bulk compression. As a proof-of-concept, the NatureLM is prompted to generate materials with the following instructions:
@@ 5.3 Material generation for desired properties
We evaluated the SMACT validity, stability, novelty, and precision of the generated materials. Success rate is defined as the ratio of generated materials whose bulk modulus is within 10% of the instructed value, compared to other generated materials.
@@ 6.2 Guide RNA design
Guide RNA (gRNA) plays a critical role in CRISPR-Cas9 gene editing by directing the Cas9 nuclease to a specific genomic target site. A gRNA consists of two components: a crispr RNA (crRNA), which is approximately 20 nucleotides long and complementary to the target DNA sequence, and a tracrRNA, which binds to Cas9. The crRNA garners more attention as it features a variable sequence critical for targeting specific DNA, whereas the tracrRNA has a more fixed role, primarily aiding in Cas9 binding. We evaluate NatureLM on two gRNA design tasks: the first is designing crRNA for a given DNA sequence, and the second is selecting the more effective crRNA from two candidates. Examples are provided below:
@@
For the first task of designing crRNA, the NatureLM’s identifies a targeted DNA region that includes the PAM sequence (Protospacer Adjacent Motif). To obtain the final crRNA, remove the ’NGG’ PAM sequence, reverse complement the DNA sequence, and apply the central dogma principle by converting DNA into RNA (replacing all ’T’ with ’U’). These quick and easy steps ensure the resulting crRNA is ready for use. A valid crRNA must meet the following criteria: (1) Be 17 to 24 nucleotides long. (2) Match a specific region in the provided DNA sequence. (3) Be followed by an ”NGG” Protospacer Adjacent Motif (PAM) in the DNA sequence, where ”N” represents any of the four standard nucleotide bases: adenine (A), cytosine (C), guanine (G), or thymine (T).
The example demonstrates that NatureLM accurately identifies the target sequence for guide RNA, ensuring the resulting guide RNA meets all necessary criteria–with an impressive accuracy of 95.7% in designing valid gRNAs.
@@ Appendix A Supplementary figures
Figure S1: Correct ratio of property-to-molecule generation. We treat the generated molecule as a correct one if $\lvert v^{\prime}-v\lvert\leq\delta$, where $v^{\prime}$ is its property value and $v$ is the input value. Tolerance threshold $\delta$ is set to 0 for HBA, HBD, RotBonds, 0.05 for QED and FSP3, and 5 for TPSA.